### v2

Abstract
This position paper argues that the majority of theory of mind benchmarks are broken because of their inability to directly test how large language models (LLMs) adapt to new partners. This problem stems from the fact that theory of mind benchmarks for LLMs are overwhelmingly inspired by the methods used to test theory of mind in humans and fall victim to a fallacy of attributing human-like qualities to AI agents. We expect that humans will engage in a consistent reasoning process across various questions about a situation, but this is known to not be the case for current LLMs. Most theory of mind benchmarks only measure what we call literal theory of mind: the ability to predict the behavior of others. Measuring this kind of reasoning is very informative in testing the ability of agents with self-consistent reasoning. However, it is important to note the distinction between this and what we actually care about when this self-consistency cannot be taken for granted. We call this functional theory of mind: the ability to adapt to agents in-context following a rational response to predictions about their behavior. We find that top performing open source LLMs may display strong capabilities in literal theory of mind, depending on how they are prompted, but seem to struggle with functional theory of mind – even when partner policies are exceedingly simple. Simply put, strong literal theory of mind performance does not necessarily imply strong functional theory of mind performance. Achieving functional theory of mind, particularly over long interaction horizons with a partner, is a significant challenge deserving a prominent role in any meaningful LLM theory of mind evaluation.
1 Introduction
Many recent papers have been inspired by how humans are evaluated for theory of mind capabilities when evaluating theory of mind in LLMs (Bubeck et al., 2023; Kosinski, 2023; Ma et al., 2023; Street et al., 2024; Strachan et al., 2024). While this does on the surface seem like a logical course of action, it is important to remind ourselves that AI has a tendency to over optimize for its training objectives in a manner that is quite alien to the way the human brain works. For example, in Bubeck et al. (2023) they tout performance of LLMs on human theory of mind tests, but then in the very next section note a noticeable lack of ”process consistency” in LLM explanations. What this means is that LLMs can come up with compelling explanations for what they do that have very little to do with their actual reasoning process. We must therefore proceed with caution about our conclusions when we evaluate LLMs in a way that does not directly align with what we really care about. When we evaluate humans, we typically focus on what we term in this paper to be literal theory of mind, which is their ability to predict the behavior of other agents. As noted by Ma et al. (2023) this can take the form of various abstractions of this behavior including actions, intentions, beliefs, percepts, desires, knowledge, and emotions. However, with humans we take for granted that this prediction of the behavior of other agents will be consistently applied into their own reasoning process when determining their own behavior. The main insight of our paper is that this process consistency cannot be taken for granted when evaluating LLMs.
To be concrete, our position is that: Current LLM theory of mind benchmarks are broken because they only measure literal theory of mind (Definition 2.1) rather than the functional theory of mind (Definition 2.2) that we actually care about our agents displaying in most cases. In Section 4 we demonstrate simple concrete cases where top open source LLMs do display strong literal theory of mind performance that is illusory in nature because it does not also lead to strong functional theory of mind performance. This serves as a case study to highlight the misleading nature of most LLM theory of mind benchmarks even when we do find that agents achieve impressive performance.
Recently, generative AI, particularly in the form of LLM assistants, has been deployed as a tool for a growing variety of real-world use cases where the LLM must interact with a diverse set of people performing a diverse set of tasks. As typically deployed today, these LLMs are only interacting with users at inference time due to the significant computational cost of continuing to train these large models in the context of individual users. As a result, LLMs must learn to adapt their behavior to users in-context based on recorded interaction histories. Indeed, a similar scenario of growing importance involves LLMs interacting with other AI agents through the agentic framework (Shavit et al., 2023). Our work is inspired by the work of Akata et al. (2023) to leverage canonical repeated games from behavioral game theory as a way to asses the adaptation ability of LLMs across the full spectrum of incentive structures. Unfortunately, we find significant deficiencies in the ability of top open source LLMs to adapt to new scenarios in-context, providing a sobering analysis of the ability of LLMs to reliably adapt without continual training of the model weights themselves.
The position of our paper was inspired by an unexpected result in our early experiments. We considered the incredibly simple scenario of an LLM agent playing the classic game Rock, Paper, Scissors against an agent that always plays the action ”Rock” for 100 consecutive rounds. The optimal course of action is for the LLM to respond with the counter to this action ”Paper” as much as possible. If you were in the LLMs shoes how do you think you would do? It is certainly reasonable to take a few exploratory actions, but there would be little excuse for not winning a high percentage of the 100 rounds. What we found is that a vanilla application of common open source LLMs resulted in a policy that chose each action ”Rock”, ”Paper”, and ”Scissor” roughly evenly. This is interesting because this is actually the famous Nash equilibrium solution for this game. However, this is the solution for optimizing your worst case return across any possible opponent. To act in this way against this particular opponent that always plays ”Rock” for 100 consecutive rounds actually demonstrates a profound lack of theory of mind. But it should be emphasized that this LLM agent does actually display a high-level of literal theory of mind. Indeed, it only takes so many rounds of seeing the other agent take the ”Rock” action before the LLM is able to predict it will keep doing so. The problem is that this prediction has nothing to do with its own chosen course of behavior. This is akin to a chatbot that continues to provide a certain response type to a user even when it would predict this response will lead to a reaction where the user expresses that they are displeased with it. We believe this is an important insight for the AI community to grapple with when measuring theory of mind in agents, and formalize the concept of functional theory of mind to accurately describe the lacking capability.
2 Two Types of Theory of Mind Between Multi-Agent LLM Policies
The Environment. The interaction process between multiple agents is often formalized as a stochastic game (Shapley, 1953) or rather a Markov game in the fully observable setting (Littman, 1994). For generality in this work, we rather consider the setting of partially observable stochastic games (POSGs) (Kuhn, 1953) that generalize POMDPs (Kaelbling et al., 1998) and MDPs (Puterman, 1994) to the multi-agent and decentralized setting. A POSG can be defined as the tuple . Here is a finite set of agents, is a finite set of global states, and is the set of joint actions across agents . is the state transition function based on the joint actions across agents, and is the joint reward function with reward function for each agent . is the joint set of observations with denoting a finite set of observations for each agent. is the function that produces agent observations based on the state. Finally, is the horizon of interactions before termination.
Agent Interaction. We can now consider environment interaction at each step from the perspective of a given focal agent where we will use to denote set of all other agents. At step agent takes action and the other agents take a joint action yielding a transition from to with probability where . The agents then receives its own observation of this state and reward . In the environments we consider in our paper it is assumed that also contains information about the actions of the other agents at the previous step . However, this need not always be the case. In such a POMDP, each agent’s policy generates its action stochastically based on its own interaction history where . Note that policies being defined in this way subsumes the common case where part of the history is discarded for computational or memory efficiency.
LLM History Representations. As we are interested in evaluating LLMs for this problem it is additionally assumed that this history representation must be encoded in text (i.e. LLM tokens). This implies that there must be a function that we have direct access to converting the observations , actions , and rewards to the form of LLM token representations. Moreover, it is important to note that we are interested in learning generalist policies across tasks with LLMs. As a result, following the protocol of Akata et al. (2023), it can also be useful to include any information about the environment tuple to the agent to promote rapid adaptation to new tasks.
2.1 Disentangling Literal Theory of Mind from Functional Theory of Mind Evaluation
The goal of any individual agent interacting in an environment for steps is to learn a policy that maximizes its expected reward given the policies of the other agents in the environment starting from state :
Because the best course of action is critically dependent on the policies of the other agents , it is common for agents to directly learn a model of the other agent’s policies in decentralized settings. This is analogous to model-based RL (Sutton & Barto, 2018; Brafman & Tennenholtz, 2002; Schrittwieser et al., 2020) and is useful for stabilizing learning (Lowe et al., 2017). However, it is worth noting that it is not a strict requirement for representing a policy that is optimal in functionality (Sutton, 1991). That said, it is tempting to consider when using LLMs as it can be generated by simply prompting the model with a token representation of the interaction history for . We can then model performance of any particular approximate literal theory of mind model using Definition 2.1.
[backgroundcolor=blue!5!white,linecolor=blue!75!black,linewidth=1pt]
Definition 2.1 (-Step Literal Theory of Mind Loss).
The loss from start state with respect to a joint policy that generates actions of its approximation that generates actions is:
where is some distance function and is some abstraction mapping function over actions.
We phrase Definition 2.1 in terms of an abstract distance function and abstraction mapping function to keep the definition as broad as possible and encompass as much as we can of the existing literature on theory of mind. Primitive actions can definitely be directly considered (as we do in this paper). In this case, is the identity mapping and is set to a percent error metric. can also be straightforwardly set to various temporal abstractions of actions given sufficient history length . Moreover, latent features that impact the behavior of agents such as intentions, beliefs, percepts, desires, knowledge, and emotions can also be considered as alternative for – as an inverse mapping must exist. However, the agent’s actions alone may not provide enough information, for example, to fully determine an agent’s emotional state, we can only judge the theory of mind performance of an agent in terms of the information provided. This does not serve as meaningful limitation of this definition because Bayesian reasoning is generally the desired outcome in the presence of lack of information.
One obvious issue with Definition 2.1 is that it is a function of a theory of mind model , which is not even necessarily needed to express the policy of an agent . For cases where the value of is decoupled from the reasoning involved in , it is of particular importance to consider a functional metric of the degree to which the policy is catered to the particular other agent policies . We can define this metric in terms of the -step regret incurred by .
[backgroundcolor=blue!5!white,linecolor=blue!75!black,linewidth=1pt]
Definition 2.2 (T-Step Functional Theory of Mind Regret).
The step regret from start state of policy that receives individual rewards when playing with policy in comparison to the optimal policy that plays the optimal -step response to policy and receives rewards :
Definition 2.2 provides us with a functional metric for measuring theory of mind in the presence of other agents parameterized by . However, it is still worth considering when the conclusions from this metric will be much different than conclusions from a literal theory of mind metric following Definition 2.1. To do this we must define a new policy that is directly based on the literal theory of mind model prediction such that behavior is chosen to be a rational maximization of the action-value function. In our experiments we simulate such a policy utilizing ground truth knowledge of the payoff structure and use to denote the regret of such a policy. thus represents the best regret that can be possibly achieved by faithfully following the predictions of our theory of mind model under the assumption that the predictions are correct.
Metrics for our Experiments. In our experiments, we aim to get a holistic view of both the literal theory of mind and functional theory of mind performance of each LLM and prompting strategy. We report the accuracy of the literal theory of mind predictions with respect to individual actions as ToM %. We also report the regret per step functionally achieved by each policy as . Finally, to get a clearer picture of the difference between the literal theory of mind performance and functional theory of mind performance, we report the regret per step of the rational policy based on the literal theory of mind model.
Accumulated Reward vs. Regret. In the experiments of our paper, the games are simple enough that it is easy to compute the reward rate of the optimal policy. However, in most cases with complex interactions the optimal policy is unknown or even unknowable. In these cases, the accumulated reward is commonly used in place of regret. This is also a valid metric of functional theory of mind as it only differs from the regret by a constant factor based on the accumulated reward of the optimal policy. The advantage of using regret, when possible, is to directly keep track of the relative optimality.
2.2 Initial Results
LLM Model | ToM % | ||
Tabular | 0.083 0.004 | 0.039 0.006 | 97.4 0.4 |
LLAMA-2 70B Chat | 0.857 0.142 | 0.119 0.006 | 92.1 0.4 |
LLAMA-2 70B | 0.971 0.067 | 0.048 0.003 | 96.8 0.2 |
LLAMA-2 13B Chat | 0.891 0.173 | 0.095 0.006 | 93.7 0.4 |
LLAMA-2 13B | 1.015 0.044 | 0.049 0.003 | 96.7 0.2 |
LLAMA-2 7B Chat | 0.904 0.151 | 0.085 0.004 | 94.3 0.3 |
LLAMA-2 7B | 0.972 0.039 | 0.066 0.004 | 95.6 0.3 |
Falcon 40B | 0.973 0.040 | 0.056 0.005 | 96.2 0.3 |
Mixtral 8x7B Instruct v1 | 0.542 0.070 | 0.050 0.005 | 96.7 0.3 |
We begin by conducting experiments on the Rock, Paper, Scissor domain as discussed in the Introduction. LLAMA-2 had previously been found to achieve strong performance on repeated matrix games (Lorè & Heydari, 2023), so we considered initial experiments with the full LLAMA-2 family (Touvron et al., 2023) as well as the competitive Falcon 40B (Almazrouei et al., 2023) and Mixtral models (Jiang et al., 2024). Our results are provided in Table 1. Our prompting strategy for the LLM policy follows what was established for repeated matrix games by Akata et al. (2023) (see Figure 1). Our prompting strategy for the theory of mind model is detailed in Figure 2. The horizon of interaction is steps against a randomly chosen fixed single action policy i.e. always Rock, always Paper, or always Scissors. 95% confidence intervals of the mean estimate across opponent policies are provided. See Appendix B for further details.
We also implemented a simple tabular model that performs efficient exploration following the classic RMax algorithm (Brafman & Tennenholtz, 2002) for near optimal worst case sample efficiency driven by optimistic Q-Learning (Watkins, 1989) with the actions at the last time step treated as the state representation. The tabular theory of mind model is based on a simple frequency count based prediction using the same state representation. It is clear that no LLM model gets even close to the performance of the tabular model in terms of functional theory of mind performance – although some come close in terms of their literal theory of mind performance. We find that Mixtral outperforms the LLAMA-2 family and we will thus subsequently include experiments with LLAMA-3 (which performs better). However, it is also interesting to see the trends within the LLAMA-2 family of models. Bigger models seem to get better performance and models with instruction tuning (labeled as “Chat” models) tend to get worse literal theory of mind performance, but better functional theory of mind performance. This is an intuitive result as instruction tuning further prioritizes interactive decision making and takes the model further from its original language modeling objective that is key to predicting what will come next in a sequence of interactions.
3 Overview of Existing Benchmarks
Frameworks for Machine Theory of Mind. In the field of multi-agent reinforcement learning, theory of mind is generally interpreted as the ability to directly predict the behavior (i.e. the actions) that another agent will take (Lowe et al., 2017; Rabinowitz et al., 2018). However, there can be a number of different ways to abstractly represent this behavior (Ma et al., 2023). For example, abstraction can be applied in the temporal dimension representing behavior as a composition of hierarchical skills (Sutton et al., 1999; Bacon et al., 2017; Riemer et al., 2018b; 2020; Allen et al., 2020; Abdulhai et al., 2021) or goals (Schaul et al., 2015; Andrychowicz et al., 2017; Yang et al., 2018; Kim et al., 2019a; b). Abstraction can also be applied at the agent level in order to represent the abstract actions of groups (Memarian et al., 2022; Touzel et al., 2024). Moreover, as showcased by the games Hanabi (Bard et al., 2020; Nekoei et al., 2021; Malloy et al., 2021a; Nekoei et al., 2023) and Poker (Ganzfried & Sandholm, 2015; Brown & Sandholm, 2019), theory of mind can be directly related to inferring recursive beliefs about unobserved information or knowledge (Moreno et al., 2021). We consider the efficacy of predictions of all of these representations of behavior as examples of measuring literal theory of mind (Definition 2.1). However, it has been recently recognized in the human-computer interaction (HCI) literature that we must go even further to achieve mutual theory of mind for better collaboration with humans (Wang & Goel, 2022; Zhang et al., 2024). In this framework, the mutual shaping of mental representations as well as the functional goal of improved performance on tasks are emphasized. We are inspired by this aspirational goal in our paper, leading us to define a functional measurement of theory of mind performance (Definition 2.2).
Measuring LLM Theory of Mind. While a number of recent studies have touted superior human-level LLM theory of mind capabilities, to the best of our knowledge these tasks have always center around passive question answering (Bubeck et al., 2023; Strachan et al., 2024; Kosinski, 2023; Street et al., 2024) as typified by the classic Sally-Anne false-belief test (Baron-Cohen et al., 1985). LLMs have also been successfully employed to simulate a diversity of personas (Park et al., 2024; Ha et al., 2024), trust behaviors (Xie et al., 2024), or even for macro-economic simulations (Li et al., 2023). However, these tasks all lack interactivity and only reflect strong performance at literal theory of mind (Definition 2.1). On the other hand, Kim et al. (2023) found that LLMs perform poorly when subjected to multiple question types that demand the same consistent underlying reasoning. This finding is in the same spirit and complimentary to the finding of our paper as well as the finding that LLMs lack process consistency (Bubeck et al., 2023). Moreover, LLMs have been found to fail at important interactive applications such as adaptive eduction for users of diverse age or education levels (Rooein et al., 2023) and providing coding assistance for beginner programmers (Nguyen et al., 2024).
Behavioral Game Theory with LLMs. Our analysis of functional theory of mind in LLMs closely follows work considering LLMs in the context of behavioral game theory. However, this work has not featured contextualization of literal theory of mind performance and is not included as an aspect of popular benchmarks focusing on measuring literal theory of mind. Our paper builds off the work of Akata et al. (2023) who considered LLMs playing the Prisoner’s Dilemma and Battle of Sexes. However, we focus on performance playing with simple policies rather than playing with other LLMs. LLMs will generally tend to generate in-distribution interaction histories when they play themselves, so we also find it more interesting to focus on very simple relatively out of distribution partner policies. For the Battle of Sexes game, Akata et al. (2023) found that GPT-4 only could adapt the canonical human-like alternating policy when using a special form of social cognition prompting that considers its own predictions of the actions of others (similar to ”perspective taking” prompting (Xu et al., 2024)). We also tried this approach in our setting and did not find it to be an adequate fix the issues that LLMs experience. The ability for LLMs to play Rock, Paper, Scissors was previously considered by Fan et al. (2024) and the ability for LLMs to play the Prisoner’s Dilemma was previously considered by Lorè & Heydari (2023), but not in the context of theory of mind.
4 Further Analysis: Difficulties in Achieving Functional Theory of Mind
Prompting | Game | ToM % | ||
LM | RPS | 0.542 0.070 | 0.050 0.005 | 96.7 0.3 |
QA | RPS | 0.881 0.117 | 0.606 0.043 | 59.6 2.9 |
CoT | RPS | 0.648 0.042 | 0.998 0.020 | 33.5 1.4 |
Plans + Insights | RPS | 0.705 0.058 | 0.443 0.041 | 70.5 2.7 |
Reflexion | RPS | 0.759 0.033 | 0.530 0.042 | 64.7 2.8 |
Social QA | RPS | 0.676 0.058 | 0.619 0.045 | 58.7 3.0 |
Social LM | RPS | 0.643 0.058 | 0.119 0.038 | 92.0 2.5 |
LM | IBS | 2.055 0.391 | 0.216 0.026 | 97.3 0.4 |
QA | IBS | 2.518 0.422 | 1.818 0.259 | 75.9 4.0 |
CoT | IBS | 2.169 0.107 | 1.659 0.095 | 79.8 1.4 |
Plans + Insights | IBS | 2.806 0.279 | 1.800 0.274 | 75.5 4.1 |
Reflexion | IBS | 2.590 0.179 | 1.726 0.270 | 77.9 4.2 |
Social QA | IBS | 2.557 0.451 | 2.190 0.383 | 70.3 5.7 |
Social LM | IBS | 2.082 0.329 | 0.182 0.057 | 97.4 0.8 |
LM | IPD | 0.949 0.142 | 0.098 0.029 | 97.8 0.5 |
QA | IPD | 2.365 0.179 | 1.342 0.177 | 68.8 2.7 |
CoT | IPD | 0.955 0.115 | 1.105 0.041 | 60.6 4.5 |
Plans + Insights | IPD | 1.379 0.180 | 1.169 0.195 | 72.3 3.2 |
Reflexion | IPD | 1.403 0.156 | 1.207 0.209 | 69.5 2.9 |
Social QA | IPD | 1.569 0.132 | 1.338 0.234 | 68.2 3.9 |
Social LM | IPD | 1.098 0.089 | 0.110 0.037 | 97.4 0.8 |
In this section, we attempt to gain a better understanding of the surprising result in Section 2.2. Rocks, Paper, Scissors (RPS) is a canonical competitive game – as can be seen by the payoff table in Figure 4. We also wanted to evaluate our models on a canonical cooperative game, for which we chose the Iterated Battle of Sexes (IBS) following Akata et al. (2023) with payoffs detailed in Figure 5. Once again following Akata et al. (2023), we also evaluate our models on the famous Iterated Prisoner’s Dilemma (IPD) mixed incentive game with payoffs detailed in Figure 6. We opted to test each model on the full spectrum of incentive structures to rule out explanations for performance lacking in competitive games such as an intrinsic altruism bias in LLMs (as has been previously suggested by Leng & Yuan (2023)).
4.1 Different Prompting Strategies
We consider three high-level classes of prompting strategies in order to provide a comprehensive evaluation.
Generic Strategies: We will henceforth call the prompting strategy used in Section 2.2 LM Prompting (see Figure 1) where the probability of each action in the action space is explicitly drawn from the model to generate the next action via next token prediction. See Figure 2 for the literal theory of mind version of the prompt. In QA Prompting (Figures 4 and 5), decision making is posed as a question answering problem where the LLM keeps generating stochastic outputs until they are in the desired format and action space vocabulary. This is implemented as a step along the way to Chain of Thought (CoT) reasoning (Wei et al., 2022) that we refer to as CoT Prompting (Figures 6 and 7) in which the LLM must generate a reasoning process and its own answer based on that reasoning process in the correct format.
Strategies for In-Context RL: We also wanted to consider prompting strategies in the literature specifically targeted at adapting LLMs in context to maximize reward. Fish et al. (2024) introduced Plans + Insights Prompting as a way to promote coordination between LLM agents based on GPT-4 by having the agents also generate text related to plans and insights and not just actions at each step while feeding in these generated plans and insights an input at the next step (see Figures 8 and 9). The idea is to be helpful in promoting consistent strategies across LLM calls. We also consider Reflexion Prompting (Shinn et al., 2024), which is a method for performing so-called verbal RL in-context by keeping a memory of reflections updated each time feedback is received from the environment. We provide example prompts in Figures 10, 11, and 12 and consider memory sizes between 1 and 3 (we always report the best value).
Strategies Explicitly Considering Theory of Mind: Finally, we consider Social Prompting as proposed for IBS by Akata et al. (2023) where the LLM first generates a prediction of the other agent’s action and then conditions its reasoning based on that action. This approach is detailed in Figure 18 with the variant using QA prompting for the action prediction (Figure 5) being called Social QA Prompting and the one using LM Prompting for the action prediction (Figure 3) being called Social LM Prompting.
Our results applying these strategies to Mixtral 8x7B Instruct v1 are included in Table 2. We also provide comprehensive results in Appendix B.2 including for the LLAMA-2 70B Chat and LLAMA-3 70B Instruct models (Table 7) and Mistral Large 2 (Table 8). We find that LM Prompting always achieves the best literal theory of mind performance. Adding the QA format seems to make literal theory of mind a lot worse by taking the input distribution further from the LLM task and closer to the kind of processing needed for actual decision making. The gap between functional theory of mind and literal theory of mind performance is predictably smaller in this case. CoT Prompting has an inconsistent effect on literal theory of mind and always appears to make the functional theory of mind performance better than vanilla QA Prompting. Reflexion and Plans+Insights also seem to improve on QA prompting in terms of functional theory of mind performance, although we find that they do not quite match the performance of generic CoT prompting.
Perhaps the most interesting case is Social Prompting, which seems consistently helpful regardless of the method of literal theory of mind prompting. However, a large gap between functional theory of mind and literal theory of mind remains even with Social Prompting. This is counterintuitive because the literal theory of mind model is directly provided as an input. It appears that the LLM still does not generate rational responses to these predicted actions. Note that the discrepancy in the LM Prompting ToM % comes from the need to prompt before agent’s action is generated for Social Prompting (Figure 3) while prompting after the action is generated (Figure 2) is more effective when possible.
4.2 Reasoning Over Long Contexts
Due to the surprising result of Social Prompting still experiencing a significant functional theory of mind gap, we wanted to understand more about the difficulty these LLMs may experience when reasoning over a long context. We now consider the LLAMA-3 70B Instruct model (Dubey et al., 2024) for its superior performance over long context tasks. We also add Oracle Prompting where the actual action the partner will take (and not just a prediction) is directly provided as input (Figure 15). As well as Oracle + Max Prompting where maximizing the reward in response to their action is further emphasized. We additionally consider a variant where the interaction history or the payoffs are removed from the prompt. Finally, we added a variant of CoT Prompting where three in-context examples of ideal thought processes are provided (Figure 13) that we call CoT 3-Shot.
Prompting | RPS | IBS | IPD |
Tabular | 0.083 0.004 | 0.211 0.012 | 0.086 0.009 |
QA | 0.444 0.107 | 1.391 0.283 | 0.996 0.133 |
CoT | 0.213 0.015 | 2.475 0.208 | 0.892 0.210 |
CoT + 3-Shot | 0.121 0.017 | 0.526 0.081 | 2.773 0.730 |
S2A | 0.224 0.027 | 1.855 0.314 | 0.808 0.192 |
S2A + CoT | 0.234 0.014 | 2.030 0.287 | 0.679 0.137 |
Social QA | 0.256 0.017 | 1.613 0.163 | 0.550 0.071 |
Social LM | 0.378 0.052 | 3.437 0.154 | 0.803 0.097 |
Plans + Insights | 0.173 0.026 | 2.405 0.216 | 0.711 0.114 |
Reflexion | 0.465 0.021 | 3.348 0.105 | 1.076 0.139 |
Oracle | 0.238 0.046 | 1.343 0.146 | 0.839 0.056 |
Oracle + Max | 0.153 0.014 | 0.785 0.072 | 0.767 0.100 |
Oracle +Max -History | 1.275 0.152 | 2.060 0.079 | 0.206 0.018 |
Oracle +Max -Payoffs | 0.103 0.015 | 0.883 0.097 | 0.241 0.046 |
In Table 3 we provide the regret per step of each prompting strategy across the three games. LLAMA-3 seems to consistently outperform LLAMA-2, but even still Social Prompting does not close the gap with the tabular model. Meanwhile, CoT leads to big improvements sometimes, but is inconsistent, making performance worse for same cases.
Indeed, it is surprising to see that Oracle consistently performs worse than the tabular RMax model that is learned from scratch without access to the payoff table or knowledge of the other agent’s policy. This speaks to a difficulty reasoning over the long contexts of this task. Incentivising maximization in the prompt definitely improves performance, but does not change the overall picture.
For RPS and IBS the interaction history seems vital for performance and the payoff table is not, implying the LLM must strugle to effectively reason about the payoffs. For IPD the combo of both the payoff table and interaction history is destructive such that it does much better with either in isolation. Seeing that the difficulty of reasoning over long contexts is at the core of the issue, we also tried a popular approach called System 2 Attention (S2A) (Weston & Sukhbaatar, 2023) to summarize the payoffs and history before sending it to either QA or CoT Prompting (Figures 16 and 17). S2A adds some value to CoT for IPD, but it is not as good as CoT for RPS and IBS. Still no LLM matches tabular performance.
Deeper Analysis. We have conducted a number of experiments that we do not have space to recap in detail within the main text. In Appendix B.3 we validate the generality of our findings regarding functional theory of mind performance across open source models including LLAMA-3 70B Instruct, Mixtral 8x7B Instruct v1, and Mistral Large 2 when paired with more dynamic tit-for-tat style partner policies. Additionally, in Appendix B.4 we look at the role of inductive bias in the action representation on performance of LLAMA-3 70B Instruct and Mistral Large 2 models. We find that more inductive bias is helpful with a small number of interactions, but could lead to poor long-term behavior that does not converge with respect to the partner’s policy.
5 Alternative Views
In this section, we take some time to discuss and refute important alternative views to the position of our paper.
We should embrace game theory, not theory of mind. This is perhaps the most prominent alternative view to our paper that was already highlighted in our Rock, Paper, Scissors example. Many papers in the multi-agent RL literature consider finding a game theoretic equilibrium solution as their ultimate goal (Littman et al., 2001; Wang & Sandholm, 2002; Greenwald et al., 2003; Zinkevich et al., 2005; 2007). In some ways game theoretic solutions are opposite to solutions that are found with theory of mind because they focus on performance against worst case or optimal partners rather than the partners agents they are actually paired with. Likewise, they focus on eliminating the exploitability of solutions by other agents as opposed to maximizing expected return in the context of the actual policies of these agents. Proponents of the game theoretic approach will readily acknowledge that theory of mind is more descriptive of the way that humans tend to act (Colman, 2003; Larson, 2004; Owen, 2013). The argument is rather that game theory is prescriptive of the way that AI agents should ideally act. We believe that there is indeed some validity to this point and acknowledge that in a closed system in which all involved agents are AI, this is the optimal case. However, this does not hold as the community moves towards increasingly real-world use cases in which AI must interact with humans or AI agents that behave sub-optimally. These use cases are the norm for LLMs and are the setting in which researchers are interested in measuring LLM theory of mind capabilities. Here there is a gap between the expected reward achievable with the theory of mind solution that optimizes for the expected case and the game theoretic solution that optimizes for the worst case. Moreover, this gap should only get larger as the agent interacts with more humans. We definitely acknowledge that what society will end up preferring depends on the alignment of incentive structures. Humans will probably prefer theory of mind solutions when their relationship with the agent is mostly collaborative, but will not want to be exploited when their relationship with the agent is mostly adversarial. We strongly support instituting guardrails in these cases and believe that this makes it even more important to measure functional theory of mind performance to understand the extent of exploitation by AI models. We provide further detail about the connection between functional theory of mind and relevant equilibria concepts in Appendix A.
Literal theory of mind is paramount for insights related applications. For some applications, we only desire that AI agents provide us with insights, which we will use to make decisions. For these applications, we do not wish to provide the AI with the agency to make these decisions for us. There are indeed very important use cases of this type for which literal theory of mind would be a sufficient capability without functional theory of mind. For example, providing theory of mind insights to humans that struggle with theory of mind (such as those with autism (Baron-Cohen, 1989; Perner et al., 1989; Tager-Flusberg, 1992; Fletcher-Watson et al., 2014)) or providing insights about the future behavior of foreign governments (Ye et al., 2024). Generally, these applications are most helpful when the AI has access to more information than the human pertaining to the situation. However, often times the amount of agency we decide to give to an AI depends on its expected reliability in making actual decisions. In this paper we argue that functional theory of mind serves as a necessary test for deciding if AI is ready to receive this agency and that literal theory of mind on its own is not reliable for assessing this performance. It is also important to note that in most applications where we do not provide AI with the agency to act, it would still be useful to humans to have the AI provide an action recommendation.
The word ”broken” is too strong. It can be argued that because there are many ways to measure literal theory of mind, adding functional theory of mind is a just single additional aspect within a comprehensive theory of mind evaluation. However, we think ”broken” is an appropriate term here because functional theory of mind is what we really care about for applications where we give the AI agency to act, and we have found that deploying an LLM in this setting simply based on its literal theory of mind performance can be misleading to a potentially dangerous extent depending on the application. It has also been argued that process consistency will simply emerge with larger and more capable models (Bubeck et al., 2023). The position of our paper is agnostic to the method by which process consistency is achieved. Regardless, given that models like GPT-4 do not display process consistency (Bubeck et al., 2023), we believe it is of paramount importance to include measurements of functional theory of mind to directly measure the behavioral implications of the achieved level of process consistency in the context of assessing theory of mind performance.
Functional theory of mind is easy with fine-tuning. RL style fine-tuning would easily solve the functional theory of mind problem displayed in our matrix game based experiments as highlighted with simple tabular models. However, our experiments still demonstrate a case where LLM literal theory of mind capabilities do not imply functional theory of mind capabilities and thus highlight the importance of directly measuring functional theory of mind. Most prominent LLM based services today also do not actually tune the model weights separately for each user and rely on in-context learning. The likely reason for this is partly about hardware related restrictions and partly about the difficulty of performing reliable continual learning. See Appendix A for a detailed discussion of connections to continual learning.
6 Conclusion
In this paper, we have argued that existing theory of mind benchmarks are broken because they measure literal theory of mind (Definition 2.1), and not functional theory of mind (Definition 2.2), which is what we typically actually care about in multi-agent social contexts. In Section 4 and Appendix B we provided comprehensive evidence that functional theory of mind performance can be quite poor even when literal theory of mind performance is good. Our analysis considers LLMs playing simple matrix games with very simple partners to underscore this point. This position paper serves as a call to action to the community to develop suitably complex LLM theory of mind benchmarks that directly measure functional theory of mind performance when acting with partners from different personas and social contexts.
Impact Statement
Our work aims to improve the theory of mind capabilities of AI agents by arguing that those capabilities depend both on making accurate predictions and acting upon those predictions in a rational way. If our viewpoint is widely adopted, AI agents will materially improve in their ability to respond to human intention and behavior. At a societal level, these improvements would be beneficial, such as through improved productivity from having more efficient AI agent user experiences. However, these improvements may also be harmful if the AI’s theory of mind is used to manipulate human behavior in nefarious, malicious, or undisclosed ways.
Appendix A Connections to Equilibria Concepts, Continual Learning, and Potential Applications
Functional Theory of Mind Objective vs. Equilibria. The distinction between each agent maximizing their own objective with no control over the optimality of other agents and an equilibria is discussed in detail in (Kim et al., 2022a). In the terminology of that paper, each agent can maximize their search over self-stable periodic distributions of states and rewards, but it only can be considered an equilibria if every agent in the environment simultaneously achieves an optimal policy with respect to its context. Depending on the dynamism of the LLM policies, different classes of equilibria concepts may occur including Nash, Cyclic, Correlated, or Active variants (Kim et al., 2022b). In practice, the functional theory of mind objective with in-context learning matches that of a multi-agent meta-learning policy in order to adapt to an encountered distribution of other agents as in (Kim et al., 2021).
In-Context Learning vs. Continual Learning. In a multi-agent environment, the environment is nonstationary from the perspective of each agent if the policies of the other agents it interacts with are changing or learning over time (Littman, 1994). As a result, the setting of rapid adaptation to new agents in our paper can be considered a special case of continual RL (see Proposition 3 of Khetarpal et al. (2020)). Proposition 2 of Khetarpal et al. (2020) established that all continual RL problems can be modeled as partial observable problems with agents conditioned on the full interaction history. So, in principle, in-context learning over sufficiently long context representations with sufficiently expressive neural networks can be considered a general solution to continual learning problems. However, it remains to be seen if it is possible to see in-context behaviors resembling continual learning strategies like reservoir sampling replay buffers (Riemer et al., 2018a) or scalable memory efficient approximate buffers (Riemer et al., 2017a; 2019; Bashivan et al., 2019). It would be interesting to also consider policy changes from step to step and their degree of correspondence with older models similar to work leveraging knowledge distillation for continual learning (Li & Hoiem, 2017; Riemer et al., 2017b; Kirkpatrick et al., 2017). Intuitively, chain of thought reasoning may be beneficial because of its similarity with continual learning approaches that can select which layers to process at inference time (Rosenbaum et al., 2018; Cases et al., 2019; Chang et al., 2019; Rosenbaum et al., 2019a), see (Rosenbaum et al., 2019b) for a survey of approaches. More generally, this adaptive computation problem can be formalized within the coagent networks framework (Thomas, 2011; Kostas et al., 2020; Zini et al., 2020). That said, the compositional generalization needed for utilizing these models in practice makes achieving real-world success very challenging (Klinger et al., 2020), potentially making LLM style training a more viable strategy for learning this composition. That said, a word of caution comes from recent theoretical insights about the difficulty of efficiently evaluating models with large context lengths (Riemer et al., 2024a) acting in complex environments (Riemer et al., 2022). This could be a leading reason for the poor performance we are currently seeing with LLMs. Ultimately, there are many potentially interesting continual learning settings that should be considered for a comprehensive evaluation (Normandin et al., 2021), which we leave to future work.
Alternatives to Continual Training. While our work highlights the limitations of using in-context learning for continual learning with current LLMs, full fine-tuning or continual training can be quite expensive in the context of LLMs – especially when it must be done for each end user. However, more computationally efficient alternatives like parameter efficient tuning (Hu et al., 2021; Dettmers et al., 2024; Thakkar et al., 2024a) or model merging (Ilharco et al., 2022; Yadav et al., 2024; Akiba et al., 2024; Thakkar et al., 2024b) may constitute a more economical middle ground. We leave exploration of the comparative efficacy of these strategies to future work. It is important to note that neural scaling laws only demonstrate generalization improvements with bigger model sizes when there is access to as much data as needed (Kaplan et al., 2020). In fact, capacity limits have been found to enhance generalization for RL (Malloy et al., 2020b; 2023) and multi-agent RL (Malloy et al., 2020a; 2021b; 2021a) in the limited data regime. Smaller models also may be necessary to maintain performance in realtime environments that are sufficiently stochastic (Riemer et al., 2024c; b). As such, it may make more sense to consider smaller LLMs if we plan on doing user specific finetuning or auxiliary objectives we would like to maintain like moral values (Padhi et al., 2024; Dognin et al., 2024).
Applicability Beyond Multi-agent RL. In this paper, we focused on multi-agent RL environments because of our use of this formalism in Definitions 2.1 and 2.2. However, the principles discussed in this paper should apply to many practical domains that are not typically modeled using RL such as biomedical applications (Das et al., 2018), making decisions based on conversation topics across the internet (Riemer et al., 2015; Heath et al., 2015; Riemer, 2017; Khabiri et al., 2015), learning what information to teach other models (Omidshafiei et al., 2019), and making decisions based on incoming internet data (Riemer et al., 2016). These problems can generally be considered a special case of the RL formalism (Barto & Dietterich, 2004).
Appendix B Experiment Details
For our experiments we ran each LLM model using the Watsonx API playing a sequence of 100 step episodes with selected policies for the other agents over the course of 24 hours or a maximum of 100 episodes. Every model was evaluated for at least 30 episodes and confidence intervals are based on the actual sample size considered for each LLM model. The payoff tables used for each game are provided in Tables 4, 5, and 6. We also considered the following representations for each action as indicated in the main text:
RPS
: , , , Rock, Pasta
: , , , Paper, Rice
: , , , Scissors, Bread
IBS
: , , Fight, Pasta
: , , Ballet, Rice
IPD
: , , Cooperate, Pasta
: , , Defect, Rice
In Tables 1 and 2, the names , , and are used for , , and respectively as we only consider the RPS game. The LLAMA-2 70B Chat model is also quantized for runtime efficiency. In Tables 3 and 9 the names , , and are used for , , and respectively to promote consistency with the action spaces also used in IBS and IPD.
Partner | ||||
You | ||||
Partner | |||
You | |||
Partner | |||
You | |||
B.1 Example Prompts
In this section, we provide example prompts for each prompting type. In all cases, we tailor our provided example to the Iterated Battle of Sexes (IBS) with actions and in which the agent happens to be playing round 5 of 100. There is a slight change in terminology for Rock, Paper, Scissors (RPS) where it is said that the agents ”receive a score of +1/0/-1” rather than that they ”win 10/8/7/5/0 points”. This is just to accommodate for a negative scale of rewards in this game as we did not feel having the reward as always positive would properly reflect the incentive structure of the game to the LLMs.
![Refer to caption](extracted/6181734/images/lm_prompting.png)
![Refer to caption](extracted/6181734/images/lm_prompting_tom.png)
![Refer to caption](extracted/6181734/images/lm_prompting_tom2.png)
![Refer to caption](extracted/6181734/images/qa_prompting.png)
![Refer to caption](extracted/6181734/images/qa_prompting_tom.png)
![Refer to caption](extracted/6181734/images/cot_prompting.png)
![Refer to caption](extracted/6181734/images/cot_prompting_tom.png)
![Refer to caption](extracted/6181734/images/pi_prompting.png)
![Refer to caption](extracted/6181734/images/pi_prompting_tom.png)
![Refer to caption](extracted/6181734/images/reflection_prompting.png)
![Refer to caption](extracted/6181734/images/reflexion_prompting.png)
![Refer to caption](extracted/6181734/images/reflexion_prompting_tom.png)
![Refer to caption](extracted/6181734/images/cot_prompting_3shot.png)
![Refer to caption](extracted/6181734/images/oracle_prompting.png)
![Refer to caption](extracted/6181734/images/oracle_prompting_max.png)
![Refer to caption](extracted/6181734/images/s2a_prompting.png)
![Refer to caption](extracted/6181734/images/s2a_prompting_tom.png)
![Refer to caption](extracted/6181734/images/social_prompting.png)
B.2 Additional Details on Experiments with Single Action Partners
The best memory size for Reflexion was 1 for RPS, IBS, and IPD in Table 2. In Table 3, the best memory sizes for Reflexion was 1 for RPS and IBS, but it was 3 for IPD.
Model | Prompting | Game | ToM % | ||
LLAMA-2 70B Chat | LM | RPS | 0.857 0.142 | 0.119 0.006 | 92.1 0.4 |
LLAMA-2 70B Chat | QA | RPS | 1.123 0.236 | 0.743 0.106 | 50.5 7.1 |
LLAMA-2 70B Chat | CoT | RPS | 0.928 0.133 | 0.776 0.081 | 48.3 5.4 |
LLAMA-3 70B Instruct | QA | RPS | 0.444 0.107 | 0.414 0.027 | 72.4 1.8 |
LLAMA-3 70B Instruct | CoT | RPS | 0.213 0.015 | 0.911 0.031 | 39.3 2.1 |
LLAMA-3 70B Instruct | CoT + 3-Shot | RPS | 0.121 0.017 | 0.747 0.042 | 50.2 2.8 |
LLAMA-3 70B Instruct | S2A | RPS | 0.224 0.027 | 0.209 0.031 | 86.1 2.0 |
LLAMA-3 70B Instruct | S2A + CoT | RPS | 0.234 0.014 | 0.278 0.025 | 81.4 1.6 |
LLAMA-3 70B Instruct | Social QA | RPS | 0.256 0.017 | 0.548 0.032 | 63.5 2.1 |
LLAMA-3 70B Instruct | Social LM | RPS | 0.378 0.052 | 0.109 0.023 | 93.1 1.5 |
LLAMA-3 70B Instruct | Plans + Insights | RPS | 0.173 0.026 | 0.178 0.015 | 88.2 1.0 |
LLAMA-3 70B Instruct | Reflexion | RPS | 0.465 0.021 | 0.273 0.019 | 81.8 1.3 |
LLAMA-2 70B Chat | LM | IBS | 2.316 0.817 | 0.496 0.038 | 94.4 0.4 |
LLAMA-2 70B Chat | QA | IBS | 2.570 0.833 | 2.401 0.480 | 68.8 7.6 |
LLAMA-2 70B Chat | CoT | IBS | 3.367 0.149 | 2.301 0.120 | 71.9 2.3 |
LLAMA-3 70B Instruct | QA | IBS | 1.391 0.283 | 1.247 0.141 | 84.6 2.3 |
LLAMA-3 70B Instruct | CoT | IBS | 2.475 0.208 | 1.835 0.128 | 77.7 2.4 |
LLAMA-3 70B Instruct | CoT + 3-Shot | IBS | 0.526 0.081 | 0.997 0.106 | 88.3 1.4 |
LLAMA-3 70B Instruct | S2A | IBS | 1.855 0.314 | 1.156 0.175 | 85.3 2.9 |
LLAMA-3 70B Instruct | S2A + CoT | IBS | 2.030 0.287 | 1.300 0.104 | 83.9 1.9 |
LLAMA-3 70B Instruct | Social QA | IBS | 1.613 0.163 | 1.138 0.104 | 85.1 1.7 |
LLAMA-3 70B Instruct | Social LM | IBS | 3.437 0.154 | 0.795 0.043 | 88.9 2.4 |
LLAMA-3 70B Instruct | Plans + Insights | IBS | 2.405 0.216 | 0.756 0.097 | 90.6 1.5 |
LLAMA-3 70B Instruct | Reflexion | IBS | 3.348 0.105 | 1.117 0.127 | 85.5 2.1 |
LLAMA-2 70B Chat | LM | IPD | 2.889 0.422 | 0.220 0.034 | 93.5 0.3 |
LLAMA-2 70B Chat | QA | IPD | 3.025 0.383 | 1.665 0.430 | 61.5 7.1 |
LLAMA-2 70B Chat | CoT | IPD | 1.839 0.289 | 1.393 0.166 | 63.8 1.6 |
LLAMA-3 70B Instruct | QA | IPD | 0.996 0.133 | 0.690 0.053 | 82.3 3.0 |
LLAMA-3 70B Instruct | CoT | IPD | 0.892 0.210 | 0.834 0.056 | 72.7 3.2 |
LLAMA-3 70B Instruct | CoT + 3-Shot | IPD | 2.773 0.730 | 0.524 0.061 | 83.8 2.3 |
LLAMA-3 70B Instruct | S2A | IPD | 0.808 0.192 | 0.662 0.200 | 84.6 3.3 |
LLAMA-3 70B Instruct | S2A + CoT | IPD | 0.679 0.137 | 0.639 0.106 | 83.1 1.4 |
LLAMA-3 70B Instruct | Social QA | IPD | 0.550 0.071 | 0.638 0.036 | 78.0 2.0 |
LLAMA-3 70B Instruct | Social LM | IPD | 0.803 0.097 | 0.394 0.065 | 87.8 2.2 |
LLAMA-3 70B Instruct | Plans + Insights | IPD | 0.711 0.114 | 0.499 0.070 | 86.7 1.0 |
LLAMA-3 70B Instruct | Reflexion | IPD | 1.076 0.139 | 0.428 0.055 | 86.8 0.9 |
Model | Prompting | Game | ToM % | ||
Mistral Large 2 | QA | RPS | 0.181 0.013 | 0.133 0.01 | 91.1 0.7 |
Mistral Large 2 | S2A | RPS | 0.177 0.015 | 0.134 0.015 | 91.1 1.0 |
Mistral Large 2 | S2A + CoT | RPS | 0.127 0.012 | 0.259 0.017 | 82.8 1.1 |
Mistral Large 2 | Social QA | RPS | 0.126 0.008 | 0.081 0.006 | 94.6 0.4 |
Mistral Large 2 | Reflexion | RPS | 0.156 0.009 | 0.050 0.005 | 96.7 0.3 |
Mistral Large 2 | QA | IBS | 1.955 0.103 | 0.575 0.055 | 93.5 0.5 |
Mistral Large 2 | S2A | IBS | 0.923 0.094 | 0.497 0.051 | 94.2 0.6 |
Mistral Large 2 | S2A + CoT | IBS | 0.498 0.091 | 0.300 0.049 | 96.3 0.7 |
Mistral Large 2 | Social QA | IBS | 1.724 0.132 | 0.42 0.036 | 95.0 0.4 |
Mistral Large 2 | Reflexion | IBS | 1.076 0.058 | 0.342 0.031 | 96.0 0.4 |
Mistral Large 2 | QA | IPD | 0.688 0.045 | 0.27 0.019 | 90.4 1.0 |
Mistral Large 2 | S2A | IPD | 0.631 0.052 | 0.275 0.042 | 92.3 0.7 |
Mistral Large 2 | S2A + CoT | IPD | 0.385 0.037 | 0.255 0.027 | 91.1 1.1 |
Mistral Large 2 | Social QA | IPD | 0.585 0.062 | 0.243 0.018 | 91.2 1.1 |
Mistral Large 2 | Reflexion | IPD | 0.619 0.033 | 0.171 0.016 | 94.5 0.5 |
B.3 Additional Experiments with Simple Adaptive Partners
Our main aim is to build towards LLMs that display mutual theory of mind capabilities (Wang & Goel, 2022) in which theory of mind is used to foster coordination behavior between agents. As such, the single action agents may not be realistic as they do not shift their actions in response to the LLM agent’s actions. In Table 9 we test the prompting strategies we have considered previously, but now playing with tit for tat style strategies (Axelrod, 1980) made famous for the efficacy in IPD. Analogously, in RPS we always play the best response to the other agent’s action at the last step and in IBS we always play the same action that the other agent did at the last step. We also add the very strong Mistral Large 2 model to match LLAMA-3 models but within the Mistral family. In most cases the LLMs perform quite poorly. All models are again much worse than the tabular model with the exception of CoT Prompting for Mistral Large 2 playing IBS. In this case, it even outperforms RMax with 3 examples, which makes sense given the payoff table it knows and the examples it is given which RMax does not receive. However, this model is inconsistent and, for example, does not perform well at RPS. Interestingly, CoT prompting seems to hurt the LLAMA-3 and Mixtral models. Social Prompting also doesn’t seem to really provide benefits, which makes sense because the literal ToM performance falls off quite a bit in this setting.
Model | Prompting | RPS | IBS | IPD |
Tabular | N/A | 0.211 0.007 | 0.468 0.031 | 0.248 0.005 |
Mixtral 8x7b Instruct v1 | QA | 1.224 0.025 | 1.484 0.109 | 1.074 0.086 |
Mixtral 8x7b Instruct v1 | CoT | 1.052 0.026 | 4.654 0.222 | 2.604 0.038 |
Mixtral 8x7b Instruct v1 | CoT + 3-Shot | 1.067 0.016 | 2.590 0.190 | 1.796 0.113 |
Mixtral 8x7b Instruct v1 | S2A | 1.065 0.021 | 1.627 0.167 | 2.002 0.099 |
Mixtral 8x7b Instruct v1 | S2A + CoT | 0.994 0.026 | 3.640 0.418 | 2.675 0.051 |
Mixtral 8x7b Instruct v1 | Social LM | 1.304 0.020 | 1.257 0.100 | 1.674 0.135 |
Mixtral 8x7b Instruct v1 | Plans/Insights* | 1.538 0.023 | 2.313 0.195 | 1.791 0.112 |
Mixtral 8x7b Instruct v1 | Reflexion 1 | 1.03 0.023 | 4.169 0.255 | 2.383 0.059 |
Mixtral 8x7b Instruct v1 | Reflexion 3 | 1.076 0.015 | 3.951 0.235 | 2.213 0.064 |
Mistral Large 2 | QA | 1.064 0.018 | 2.167 0.131 | 2.581 0.048 |
Mistral Large 2 | CoT | 1.007 0.014 | 0.223 0.074 | 2.657 0.030 |
Mistral Large 2 | CoT + 3-Shot | 0.999 0.016 | 0.032 0.023 | 0.417 0.154 |
Mistral Large 2 | S2A | 1.018 0.027 | 1.009 0.129 | 0.454 0.193 |
Mistral Large 2 | S2A + CoT | 1.029 0.025 | 0.663 0.231 | 2.586 0.132 |
Mistral Large 2 | Social LM | 0.903 0.020 | 2.021 0.197 | 2.668 0.021 |
Mistral Large 2 | Plans/Insights* | 1.330 0.021 | 0.089 0.051 | 0.717 0.140 |
Mistral Large 2 | Reflexion 1 | 0.987 0.016 | 2.311 0.152 | 2.544 0.049 |
Mistral Large 2 | Reflexion 3 | 1.037 0.020 | 1.772 0.166 | 2.449 0.065 |
LLAMA-3 70B Instruct | QA | 1.174 0.023 | 1.377 0.110 | 2.770 0.041 |
LLAMA-3 70B Instruct | CoT | 1.029 0.019 | 4.468 0.176 | 2.761 0.023 |
LLAMA-3 70B Instruct | CoT + 3-Shot | 0.968 0.020 | 3.065 0.237 | 2.565 0.066 |
LLAMA-3 70B Instruct | S2A | 1.025 0.026 | 2.100 0.149 | 2.606 0.060 |
LLAMA-3 70B Instruct | S2A + CoT | 1.047 0.048 | 2.938 0.258 | 2.740 0.031 |
LLAMA-3 70B Instruct | Social LM | 1.112 0.023 | 2.448 0.335 | 2.759 0.012 |
LLAMA-3 70B Instruct | Plans/Insights* | 1.389 0.051 | 3.127 0.229 | 2.547 0.056 |
LLAMA-3 70B Instruct | Reflexion 1 | 0.953 0.023 | 5.758 0.152 | 2.57 0.031 |
LLAMA-3 70B Instruct | Reflexion 3 | 1.003 0.016 | 5.531 0.105 | 2.442 0.029 |
B.4 Additional Experiments on Action Space Inductive Bias
![Refer to caption](extracted/6181734/images/mistral_ipd_tit4tat_tiny.png)
The Good and Bad of Inductive Bias. So far we have used the neutral action representations , , and to avoid contamination following prior work (Binz & Schulz, 2023; Akata et al., 2023). However, it is interesting given our results so far to get a better sense of the degree that prior knowledge in the LLMs impacts performance and the choice of action representation is great way for us to control the extent that this knowledge is evoked. For the best performing models LLAMA-3 70B Instruct and Mistral Large 2, we conduct experiments for all three games playing with both single action and tit for tat style partners. We provide comprehensive results for functional theory of mind performance in Figures 20, 22, 24, 26, 28, 30, 32, 34, 36, 38, 40, and 42, and for literal theory of mind performance in Figures 21, 23, 25, 27, 29, 31, 33, 35, 37, 39, 41, and 43. Here we compare neutral actions with the actual canonical action names for the game, the neutral actions with 20 repetitions to take up a longer portion of the context, and nonsense word actions (see Appendix B). We generally find that both LLM models experience systematic bias that prevents them from converging to optimal performance as the number of interactions grow. Indeed, only LLAMA-3 playing IBS with tit for tat partners (Figure 28) seems on the road to slow convergence. Meanwhile, literal theory of mind seems to be converging in a greater number of settings (Figures 29, 33, 35, and 37). We find that real actions often help with functional theory of mind performance early in the interaction stream while becoming harmful for converged performance as interaction goes on (Figures 19, 20, 26, 30, 32, and 36).
![Refer to caption](extracted/6181734/images/llama3_rps_regret.png)
![Refer to caption](extracted/6181734/images/llama3_rps_acc.png)
![Refer to caption](extracted/6181734/images/llama3_ibs_regret.png)
![Refer to caption](extracted/6181734/images/llama3_ibs_acc.png)
![Refer to caption](extracted/6181734/images/llama3_ipd_regret.png)
![Refer to caption](extracted/6181734/images/llama3_ipd_acc.png)
![Refer to caption](extracted/6181734/images/llama3_rps_regret_tit4tat.png)
![Refer to caption](extracted/6181734/images/llama3_rps_acc_tit4tat.png)
![Refer to caption](extracted/6181734/images/llama3_ibs_regret_tit4tat.png)
![Refer to caption](extracted/6181734/images/llama3_ibs_acc_tit4tat.png)
![Refer to caption](extracted/6181734/images/llama3_ipd_regret_tit4tat.png)
![Refer to caption](extracted/6181734/images/llama3_ipd_acc_tit4tat.png)
![Refer to caption](extracted/6181734/images/mistral_rps_regret.png)
![Refer to caption](extracted/6181734/images/mistral_rps_acc.png)
![Refer to caption](extracted/6181734/images/mistral_ibs_regret.png)
![Refer to caption](extracted/6181734/images/mistral_ibs_acc.png)
![Refer to caption](extracted/6181734/images/mistral_ipd_regret.png)
![Refer to caption](extracted/6181734/images/mistral_ipd_acc.png)
![Refer to caption](extracted/6181734/images/mistral_rps_regret_tit4tat.png)
![Refer to caption](extracted/6181734/images/mistral_rps_acc_tit4tat.png)
![Refer to caption](extracted/6181734/images/mistral_ibs_regret_tit4tat.png)
![Refer to caption](extracted/6181734/images/mistral_ibs_acc_tit4tat.png)
![Refer to caption](extracted/6181734/images/mistral_ipd_regret_tit4tat.png)
![Refer to caption](extracted/6181734/images/mistral_ipd_acc_tit4tat.png)